Use the point-slope form to write an equation of the line with the given slope and point. Then write the equation in slope-intercept form.
Slope -4, passes through (-2, 10)

Answer by TimothyLamb(4379)

point-slope form:
y - 10 = -4(x - -2)
y - 10 = -4(x + 2)
---
slope-intercept form:
y - 10 = -4x - 8
y = -4x - 8 + 10
y = -4x + 2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php